Suppose $6,500 is divided into two bank accounts. One account pays 10% simple interest per year and the other pays 4.6%. After two years there is a total of $1000 in interest between the two accounts. How much was invested into the bank account that pays 10% simple interest (rounded to the nearest cent)?
Answer by mananth(16946)

Bank I 10.00% per annum ---x
Bank II 4.60% per annum ---y

x + 1 y= 6500 ------------------------1
10.00% x + 4.60% y= 500
Multiply by 100
10 x + 4.6 y= = 50000.00 --------2
Multiply (1) by -10
we get
-10x-10y=-65000.00
Add this to (2)
-5.4 y= = -15000
divide by -5.4
y = 2777.78 Bank I 4.60%
Balance 3722.22 Bank II 10.00%
CHECK
2777.78 --------- 4.60% ------- 127.78
3722.22 ------- 10.00% ------- 372.22
Total -------------- 500.00